OpenStudy (anonymous):

The table shows the results of a spinner with four colored sections. Use the table to answer the question. The spinner is spun 50 more times. Which is the best prediction for the number of times the spinner will stop on blue? A. 6 B. 12 C. 24 D. 48

OpenStudy (kirbykirby):

So, based on the table, you see that blue was obtained:\[\frac{24}{24+49+10+17}=\frac{24}{100}\] of the time

OpenStudy (kirbykirby):

=0.24

OpenStudy (kirbykirby):

Now, you can just calculate the fraction in the choices above (like a) is 6/50 = 0.12), and see which one matches 0.24 or is closest to 0.24, since this is the fraction of times you expect to get "blue", and this is your best guess.

OpenStudy (anonymous):

thanks answers c

OpenStudy (kirbykirby):

actually it is not C